A man travels 2 km; turn left and travel 3 km turn left again and travel 2 km how far is he from the starting point?
Answer by Fombitz(32388)

He starts at (0,0).
He moves 2 km.
(0,0)+(2,0)=(2,0)
He then moves 3 km left (which would be in the other direction.
(2,0)+(0,3)=(2,3)
When he turns left now he's moving in the original direction but in the opposite sense so,
(2,3)+(-2,0)=(0,3)
So he's 3 km from the starting position.